The coefficient of static friction between a book and the level surface it slides on is 0.65. If the mass of the book is 2.0 kg, the minimum initial applied force required to slide the books across the surface is ___N.
1
Expert's answer
2014-04-09T10:38:59-0400
F = mu*N, where mu = 0.65 is the coefficient of a static friction, N = m*g is a counterpressure
Thus, F = mu*m*g = 0.65*2.0*9.8 = 13 N.